OpenStudy (anonymous):

Factor this polynomial expression. x^3 - 216 a.(x-6)(x^2+6x+36) b.(x+6)(x^2+6x+6) c.(x+6)(x^2+6x+36) d.(x-6)(x^2+6x+6)

OpenStudy (anonymous):

Difference of cubes.

myininaya (myininaya):

\[x^3-6^3=(a-b)(a^2+ab+b^2)\] where a=x and b=6

OpenStudy (anonymous):

Answer is A ;)